OpenStudy (anonymous):

Find the value of the square root of 24 over 4 times the square root of 6?

OpenStudy (anonymous):

like this? \[\sqrt{\frac{24}{4}}*\sqrt{6}\]?

OpenStudy (anonymous):

no 24 is by its self and its 4*6 after that

OpenStudy (anonymous):

ah, i think i see. \[\frac{\sqrt{24}}{4\sqrt{6}}\]

OpenStudy (anonymous):

yeah like that

OpenStudy (anonymous):

= 2/4

OpenStudy (anonymous):

so lets simplify the square root of 24 a little: \[\sqrt{24} = \sqrt{ 6*4} = \sqrt{6}*\sqrt{4}\] Placing that on the top of fraction we obtain: \[\frac{\sqrt{24}}{4\sqrt{6}} = \frac{\sqrt{4}*\sqrt{6}}{4\sqrt{6}} = \frac{2}{4} = \frac{1}{2}\]

OpenStudy (anonymous):

because √24/√6 = √4 which is really just 2. and then the 4 is left over.

OpenStudy (anonymous):

thankss :)

OpenStudy (anonymous):

also, it works in da calcoolator.

OpenStudy (anonymous):

oh it didnt work in mine

OpenStudy (anonymous):

sha, need practice with ( )

OpenStudy (anonymous):

should of been √24/(4√6)